The linear charge density of an infinite line charge is 0.3 10 .mC - -16 x Assuming that the electric potential at a perpendicular distance of 5.0 m from the wire is zero, calculate the potential at the perpendicular distance of 6.0 m. 


1
Expert's answer
2021-06-02T09:34:26-0400

The potential at distance r from infinite charged wire is


"V=\frac{λ}{2πϵ_0}\ln(r)."

So, the potential at the perpendicular distance of 6.0 m from an infinite wire of charge "\lambda=0.3\text{ mC/m}" is


"V=\frac{0.3·10^{-3}}{2π·8.85·10^{-12}}\ln(6)=9662·10^3\text{ V}."